### v1

@@ 1.1 Background
Occasionally, the varying available memory and network bandwidth of the devices may affect the quality of the currently provided solution. Hence, model optimization algorithms should adjust their solutions accordingly to adapt to these environment changes, which means that solution quality evaluations will happen frequently. To perform the evaluations in a timely fashion due to real-time requirements, RMs (Regression Models) are commonly trained to predict relevant solution metrics - specifically in this paper, the resulted module inference latency. These trained RMs can also be applied to other adaptive optimization algorithms like inference request batching [6, 16] and input data partitioning [11, 4, 19]. Several prediction frameworks [8, 20, 10] have been introduced to specify different RM training workflows, but none of them allow flexible configurations of the input parameters (e.g., batch size, FLOPS, device utilization rate) and the of selected RMs for different DNN modules. Since different modules have varying degrees of structure and computation complexity, they require different regression models and input parameters for prediction. Hence, a flexible prediction framework that can accommodate such diversity is essential.
@@ 1.2 Gap Analysis on Related Work
In this paper, a DNN module inference latency prediction framework is proposed to address the aforementioned problem by i) hosting a set of customizable input parameters to train multiple different RMs per DNN module (e.g., convolutional layer) with self-generated datasets, and ii) automatically selecting a set of trained RMs leading to the highest possible overall prediction accuracy, while keeping the prediction time / space consumption as low as possible. Besides, a new RM named MEDN (Multi-task Encoder-Decoder Network) is proposed as an alternative solution.
@@
The rest of the paper is structured as follows: Section 1 discusses the motivation for introducing the proposed framework and briefly states its advantages compared to other existing DNN module inference latency prediction frameworks. Section 2 outlines the key components and operations of the proposed framework. Section 3 presents the experimental frame and a quantitative analysis and evaluation of the proposed system. Section 4 summarizes the paper and lists potential future research directions.
@@ 2.1 Parameters
The Sampling Parameters refer to the original parameters that can be used to construct the corresponding data and module. As an example, a convolutional module can be constructed by specifying the Sampling Parameters including the input/output channels $C_{i},C_{o}\in[3,512]$, kernel size $K\in\{1,3,5,7,9\}$, and stride $S\in\{1,2,4\}$. The corresponding input data for the convolutional module can be constructed by specifying the batch size $N\in[1,64]$, input height/weight $L\in\{224,112,56,32,28,27,14,13,8,7\}$ along with the input channels $C_{i}$. Apart from these original parameters, the available memory $M$ (in bytes) and utilization rate $U\in[0,1.0]$ of the device should be recorded as Measurable Parameters to capture the features from the dynamic environment. Intuitively, RMs have to be informed about whether the device is free and available to commit enough resources to module inference tasks. When the device is heavily occupied with other running jobs, RMs should predict a relatively slower inference speed than the normal scenario (i.e., static environment). Finally, Inferable Parameters describe the set of parameters whose values can be calculated from other Sampling/Measurable/Inferable Parameters. This category of parameters generally offers fine-grained information to help RMs better understand the input features. As an example, the number of model weight parameters $N_{m}$ indicates the module complexity and the input data size $N_{d}$ reflects the computation overheads.
@@ 3.2 Results on Single-RM Selection
According to Table 5, other RMs like RF and MLP achieve lower overall prediction accuracy and R-squared compared to the proposed MEDN. Besides, they are generally less efficient both in time (time-per-sample as "Tps" in milliseconds) and space dimension (model-size as "Size" in kilobytes). The overall metric results are essentially percentage differences with MEDN, averaged over all modules. The results prove that MEDN serves better as a alternative RM solution on all evaluated metrics.
@@ 3.3 Results on Ablation Experiment for MEDN
Figure 4 shows the evaluated metric results averaged over all modules. P20-Accuracy is the default accuracy metric in other result tables, specifying the ratio of samples whose prediction percentage error is within 20%. First, it can be noticed that MEDN outperforms MEDN-D on all metrics, verifying that the reconstruction decoder serves an important role in the improved performance. Next, MEDN-XI performs generally worse than MEDN-XM but better than MEDN-R, indicating that Inferable Parameters provide more helpful guidance to latency prediction than Measurable Parameters, although both lead to better accuracy results. Finally, it is discovered that Measurable Parameters allow MEDN to be more accurate on the correctly predicted samples, resulting in higher P10-Accuracy than that of MEDN-XM.

### v2

@@ 1.1 Background
Occasionally, the varying available memory and network bandwidth of the devices may affect the quality of the currently provided solution. Hence, model optimization algorithms should adjust their solutions accordingly to adapt to these environment changes, which means that solution quality evaluations will happen frequently. To perform the evaluations in a timely fashion due to real-time requirements, RMs (Regression Models) are commonly trained to predict relevant solution metrics - specifically in this paper, the resulted module inference latency. These trained RMs can also be applied to other adaptive optimization algorithms like inference request batching [6, 16] and input data partitioning [11, 4, 19]. Several prediction frameworks [8, 20, 10] have been introduced to specify different RM training workflows, but none of them allow flexible configurations of the input parameters (e.g., batch size, FLOPS, device utilization rate) and of the selected RMs for different DNN modules. Since different modules have varying degrees of structure and computation complexity, they require different regression models and input parameters for prediction. Hence, a flexible prediction framework that can accommodate such diversity is essential.
@@ 1.2 Gap Analysis on Related Work
In this paper, a DNN module inference latency prediction framework is proposed to address the aforementioned problem by i) hosting a set of customizable input parameters to train multiple different RMs per DNN module (e.g., convolutional layer) with self-generated datasets, and ii) automatically selecting a set of trained RMs leading to the highest possible overall prediction accuracy, while keeping the prediction time / space consumption as low as possible. Besides, a new RM named MEDN (Multi-task Encoder-Decoder Network) is proposed as an alternative solution for RM selection.
@@
The rest of the paper is structured as follows: Section 1 discusses the motivation for introducing the proposed framework and briefly states its advantages compared to other existing DNN module inference latency prediction frameworks. Section 2 outlines the key components and operations of the proposed framework. Section 3 presents the experimental frame as well as a quantitative analysis and evaluation of the proposed system. Section 4 summarizes the paper and lists potential future research directions.
@@ 2.1 Parameters
The Sampling Parameters refer to the original parameters that can be used to construct the corresponding data and module. As an example, a convolutional module can be constructed by specifying the Sampling Parameters including the input/output channels $C_{i},C_{o}\in[3,512]$, kernel size $K\in\{1,3,5,7,9\}$, and stride $S\in\{1,2,4\}$. The corresponding input data for the convolutional module can be constructed by specifying the batch size $N\in[1,64]$, input height/weight $L\in\{224,112,56,32,28,27,14,13,8,7\}$ along with the input channels $C_{i}$. Apart from these original parameters, the available memory $M$ (in bytes) and utilization rate $U\in[0,1.0]$ of the device should be recorded as Measurable Parameters to capture the features from the dynamic environment. Intuitively, RMs have to be informed about whether the device is free and available to commit enough resources to module inference tasks. When the device is heavily occupied with other running jobs, RMs should predict a relatively slower inference speed than the normal scenario (i.e., static environment). Finally, Inferable Parameters describe the set of parameters whose values can be calculated from other Sampling/Measurable/Inferable Parameters. This category of parameters generally offers fine-grained information to help RMs better understand the input features. As an example, the number of module weight parameters $N_{m}$ indicates the module complexity and the input data size $N_{d}$ reflects the computation overheads.
@@ 3.2 Results on Single-RM Selection
According to Table 5, other RMs like RF and MLP achieve lower overall prediction accuracy and R-squared compared to the proposed MEDN. Besides, they are generally less efficient both in time (time-per-sample as "Tps" in milliseconds) and space dimension (model-size as "Size" in kilobytes). The overall metric results are essentially percentage differences with MEDN, averaged over all modules. The results prove that MEDN serves better as an alternative RM solution on all evaluated metrics.
@@ 3.3 Results on Ablation Experiment for MEDN
Figure 4 shows the evaluated metric results averaged over all modules. P20-Accuracy is the default accuracy metric in other result tables, specifying the ratio of samples whose prediction percentage error is within 20%. First, it can be noticed that MEDN outperforms MEDN-D on all metrics, verifying that the reconstruction decoder serves an important role in the improved performance. Next, MEDN-XI performs generally worse than MEDN-XM but better than MEDN-R, indicating that Inferable Parameters provide more helpful guidance for latency prediction than Measurable Parameters, although both lead to better accuracy results. Finally, it is discovered that Measurable Parameters allow MEDN to be more accurate on the correctly predicted samples, resulting in higher P10-Accuracy than that of MEDN-XM.